OpenStudy (anonymous):

What is the volume of the shaded region? Round your answer to 2 decimal places if necessary. Use π = 3.14 when necessary

OpenStudy (anonymous):

This is the figure I have to use

OpenStudy (anonymous):

The volume of the cube is \[10^3=1000\], the sphere has radius 5, so it's volume is \[4/3 * \pi * r^3=4/3*3.14*5^3=523.33\] The volume of the shaded region is the volume of the cube minus the volume of the sphere so \[1000-523.33=476.67\]

OpenStudy (anonymous):

so the answer is 476.67

OpenStudy (anonymous):

476.7